A box contains 3 red and 4 green marbles. Five marbles are drawn without replacement. Let X denote the number of red marbles obtained.

a) Construct the probability distribution of X.

b) What is the expected number of red marbles?
Answer #1

a)

X                             P(X)

1                              0.143

2                              0.571

3                              0.286

b) Expected number of red marbles = n * k/N = 5 * 3/7 = 2.143